OpenStudy (anonymous):

How do I split sin(11pi/12) into sin(2pi/3 + pi/4)

OpenStudy (anonymous):

To elaborate, I mean, how did it get split into sin(2pi/3 + pi/4)

OpenStudy (anonymous):

\[\sin \left( \frac{ 11\pi }{ 12 } \right)=\sin \left( \frac{ 8\pi }{ 12 } +\frac{ 3\pi }{ 12 }\right)=\sin \left( \frac{ 2\pi }{ 3 }+\frac{\pi }{ 4 } \right)\]Get it? @sprr

OpenStudy (anonymous):

What would you split cos(7pi/8) into?